# Stable Survival Extrapolation via Transfer Learning

### v1

@@ Abstract
The mean survival is the key ingredient of the decision process in several applications, including health economic evaluations. It is defined as the area under the complete survival curve thus necessitating extrapolation of the observed data. In this article we employ a Bayesian mortality model and transfer its projections in order to construct the baseline population that acts as an anchor of the survival model. This can be seen as an implicit bias-variance trade-off in unseen data. We then propose extrapolation methods based on flexible parametric polyhazard models which can naturally accommodate diverse shapes, including non-proportional hazards and crossing survival curves while typically maintaining a natural interpretation. We estimate the mean survival and related estimands in three cases, namely breast cancer, cardiac arrhythmia and advanced melanoma. Specifically, we evaluate the survival disadvantage of triple negative breast cancer cases, the efficacy of combining immunotherapy with mRNA cancer therapeutic for advanced melanoma treatment and the suitability of implantable cardioverter defibrilators for cardiac arrhythmia. The latter is conducted in a competing risks context illustrating how working on the cause-specific hazard alone minimizes potential instability. The results suggest that the proposed approach offers a flexible, interpretable and robust approach when survival extrapolation is required.
@@ 1 Introduction
Extrapolation beyond the observed survival curve can be dangerous and should not be made casually. However, practical applications necessitate such extrapolation, especially in health economic modeling and related applications of survival analysis. This is sometimes performed by fitting naive parametric models and simply projecting them to the future. Alternatively, solutions based upon the restricted mean survival (RMS) have been proposed but from the health economic perspective these give an accurate answer to the wrong problem.
@@
Anchoring against external data has been recognised as a way to build towards robust ways of extrapolation to the future (see Rutherford et al. 2020), either from a Bayesian perspective (e.g. Demiris and Sharples 2006) or using a frequentist approach. This is essentially being proposed because it offers an implicit bias-variance trade-off in the unseen data since utilising long term information minimises the chance of wild extrapolation, albeit with the possibility of adding bias. Since this trade-off is concerned with future data, this represents a non-standard statistical problem. See Jackson et al. (2017), Bullement et al. (2023) and Latimer and Adler (2022) for recent reviews on survival extrapolation by anchoring against external data. Gallacher et al. (2021) discuss problems regarding the goodness-of-fit statistics to select a model for extrapolation while Che et al. (2023) present a method that gradually blends the predicted survival of the disease population with the survival of the general population. Guyot et al. (2017) use splines to extrapolate survival of cancer patients and Jackson (2023) provides an R package for flexible extrapolation using M-splines.
@@
Our methodology is motivated from and tested on three case studies regarding breast cancer, metastatic skin melanoma and cardiac arrhythmia. The breast cancer dataset contains long follow up and therefore facilitates the thorough assessment of the predictive ability of the proposed approach based on suitable information fractions. We use individual information on particular genes to classify individuals as triple negative or not and evaluate their life expectancy in a challenging scenario of crossing survival curves. The current developments regarding mRNA therapeutics is tackled in melanoma where we estimate the life years gained (LYG) when the mRNA vaccine is received along with pembrolizumab versus those receiving pembrolizumab immunotherapy alone. Our third use case is concerned with an arrhythmia dataset where we estimate the LYG for patients that receive an implantable cardioverter defibrilator (ICD) compared to a group receiving anti arrhythmia drugs (AAD). In doing so, we extend the poly-hazard models with the introduction of suitable change-points. The implementation of proposed methods has been carried out in the R language and the relevant Stan code is made freely available on Github.
@@ 2.2 Advanced Melanoma
Advanced melanoma refers to a type of skin cancer at metastatic stage. Their treatment has been revolutionalized due to immunotherapy such as pembrolizumab; see Villani et al. (2022) for a discussion of the currently available treatment options. Perhaps the most exciting new cancer treatment is based on mRNA technology and a number of clinical trials are currently being deployed. Khattak et al. (2023) have published the first such trial where the hazard ratio for recurrence-free survival was 0.561 for patients receiving ‘mRNA-4157 (V940) + pembro’ against pembro alone. Preliminary survival estimates of these groups are thus imperative. We digitize the published survival curve (see the on-line Appendix A) from Hamid et al. (2019) to obtain event times of patients receiving pembro. We then use the estimated hazard ratio as the basis for the ‘mRNA+pembro’ curve.
@@ 4.1 Breast Cancer
The breast cancer dataset has long follow up and hence a nearly complete survival curve, thus facilitating repeated testing of the trained models against the ‘truth’ in the form of the observed data. We use various levels of the information fraction (Demets and Lan (1994)) as a metric. For example, 80% of information corresponds to 49% survival probability. The best model for this dataset was a joint Bi-Weibull for the disease group and the external population. We impose the constraint that the first components of each Bi-Weibull model will be proportional to each other and the second components will be equal. Thus, we generate age-sex matched data from the external population $y_{1:n_{p}}$ and we assume each sample $i=1,...,n_{p}$ has hazard $h_{p}(y_{i};\alpha_{1},\alpha_{2},\lambda_{1},\lambda_{2})=h_{p}^{1}(y_{i};\alpha_{1},\lambda_{1})+h_{p}^{2}(y_{i};\alpha_{2},\lambda_{2})$. The hazard form of each component $m$ is Weibull with shape and rate parameters $\alpha_{m}$ and $\lambda_{m}$ respectively, i.e. $h_{p}^{m}(y_{i};\alpha_{m},\lambda_{m})=\lambda_{m}\cdot\alpha_{m}\cdot y_{i}^{\alpha_{m}-1}$. The survival function is $S(y_{i};\alpha_{1},\alpha_{2},\lambda_{1},\lambda_{2})=S_{p}^{1}(y_{i};\alpha_{1},\lambda_{1})\cdot S_{p}^{2}(y_{i};\alpha_{1},\lambda_{1})$, where each component takes the form $S_{p}^{m}(y_{i};\alpha_{m},\lambda_{m})=\exp(-\lambda_{m}\cdot y_{i}^{\alpha_{m}})$. The likelihood is given by equation (2). For the disease data $t_{1:n_{d}}$, we also assume a Bi-Weibull model, but every sample $i=1,...,n_{d}$ has hazard $h_{d}(t_{i};\alpha_{1},\alpha_{2},\lambda_{1},\lambda_{2})=C\cdot h_{p}^{1}(t_{i};\alpha_{1},\lambda_{1})+h_{p}^{2}(t_{i};\alpha_{2},\lambda_{2})$ and survival $S(t_{i};\alpha_{1},\alpha_{2},\lambda_{1},\lambda_{2})=S_{p}^{1}(t_{i};\alpha_{1},\lambda_{1})^{C}\cdot S_{p}^{2}(t_{i};\alpha_{1},\lambda_{1})$, where $C>0$. The likelihood is given by (1) and the complete likelihood by (3).
@@ 4.1.1 Subgroup Analysis for precision medicine
4.2 Advanced Melanoma
@@
The best fitted model for the advanced melanoma dataset was a joint Tri-Weibull where the first components of the disease and external groups are proportional to each other while the third components are equal. Thus, we use $m=3$ components: $h_{p}(y_{i};\alpha_{1},\alpha_{2},\alpha_{3},\lambda_{1},\lambda_{2},\lambda_{3})=h_{p}^{1}(y_{i};\alpha_{1},\lambda_{1})+h_{p}^{2}(y_{i};\alpha_{2},\lambda_{2})+h_{p}^{3}(y_{i};\alpha_{3},\lambda_{3})$ and the hazard of the pembrolizumab group is given by $h_{pembro}(t_{i};\alpha_{1},\alpha_{d},\alpha_{3},\lambda_{1},\lambda_{d},\lambda_{3})=C\cdot h_{p}^{1}(t_{i};\alpha_{1},\lambda_{1})+h_{d}^{2}(t_{i};\alpha_{d},\lambda_{d})+h_{p}(t_{i};\alpha_{3},\lambda_{3})$. In this example we select the ‘constant difference’ method and use the last 5 points (months) to write $h_{pembro}(t^{\prime})=D_{pembro}+h_{p}(t^{\prime})$ for a future time $t^{\prime}$. The mRNA group of patients is derived via the proportional hazards assumption $h_{mRNA}(t)=0.561\cdot h_{pembro}(t)$ (see Khattak et al. 2023) and extrapolation follows using $h_{mRNA}(t^{\prime})=D_{mRNA}+h_{p}(t^{\prime})$ for future times $t^{\prime}$.
Figure 3 depicts the extrapolated survival and hazard functions. Calculating the area between the mRNA and pembrolizumab survival functions estimates the LYG to be 43.69 months on average (see right plot of Figure 4) meaning that patients with advanced melanoma may gain 3.64 life years if the mRNA therapeutic is added to their treatment. This example also illustrates that RMS occurs early (see left plot of Figure 4) and may be an inaccurate estimate of mean survival.
@@ 5 Discussion
Fitting independent poly-hazard models to the external population and the disease groups can give seemingly good fit during the follow-up period at the cost of unstable and/or unreliable extrapolation. This article uses methods that can avoid such erratic behaviour by transferring appropriate information from data external to the disease group. The external population is based upon mortality projections as opposed to past data and the evidence is synthesized in a Bayesian context. We found that modelling on the hazard scale can identify incoherent fit more efficiently compared to the survival and cumulative hazard scale. We employ an extension of the standard poly-hazard models introducing change-points, offering added flexibility.
@@
A review of tests for crossing survival curves may be found at Li et al. (2015). Our work is concerned with a more involved issue, not simply testing for equality, but fully modelling the survival curves and the calculation of their difference, as necessitated in health economic evaluations. Alternatives to poly-hazard models include splines (Jackson 2023) and fractional polynomials (Royston and Parmar 2002). Both offer substantial flexibility at the cost of typically limited interpretability of their parameters. Perhaps more importantly, poly-hazard models offer a natural balance between efficient use of the data and undesirable carry-over effects.
@@
This work was motivated by 3 use cases that include a crossing survival problem, an emerging cancer therapeutic and a competing risks example. The unified approach presented here naturally caters for survival extrapolation in these cases. We estimate that women with breast cancer lose on average 10.17 years compared to the general population and we also estimate the additional deficit of the triple negative sub-population. The mRNA-4157 (V940) therapeutic in combination with pembrolizumab is currently being assessed for treatment of advanced melanoma and by incorporating the early trial evidence within our framework we estimate that 3.64 extra years of life may be gained compared to pembrolizumab alone. The cardiac arrhythmia estimates mortality reduction when switching from AAD to ICD treatment by retaining a common hazard on death from other causes.
@@
The results are assessed under five extrapolation methods. The baseline method performs well in all our examples by appropriately modelling the bond between the disease and external populations. The constant difference and ratio methods build upon the findings of the follow-up period and extrapolate the last observed relationship between the two populations into the future. We further amend these two methods by applying them to the components of interest alone and obtain the cause-specific analogues. The proposed methodology does require individual-level data. This problem can be resolved by digitizing published Kaplan-Meier curves as illustrated in the advanced melanoma and cardiac arrhythmia data. Still, some inaccuracies may result and care is required in this aspect. In addition, expert opinion may be elicited as done by Benaglia et al. (2015) for the age-specific proportions of arrhythmia.
@@
Survival analysis is also used in non-medical application areas of statistics, including industrial, actuarial and reliability functions where the idea of transferring information from an external source can be fruitful if extrapolation is required. Further development of the piecewise poly-hazard models may be considered by training them based upon alternative combinations of the disease and external groups. The predictive ability of the models developed in this article has been informally assessed, since in reality it is dictated by the performance on future/unseen data. One alternative is to characterize the model’s properties using an appropriate loss function and this represents the subject of current work.
@@ A. The mechanics of KM digitization
B. Some comments on usually used extrapolations
@@ C. Mortality projections
D. Breast cancer genes-specific results
Results on the extrapolation of breast cancer data related with the groups 3N and n3N.

### v2

@@ Abstract
The mean survival is the key ingredient of the decision process in several applications, notably in health economic evaluations. It is defined as the area under the complete survival curve, thus necessitating extrapolation of the observed data. This may be achieved in a more stable manner by borrowing long term evidence from registry and demographic data. Such borrowing can be seen as an implicit bias-variance trade-off in unseen data. In this article we employ a Bayesian mortality model and transfer its projections in order to construct the baseline population that acts as an anchor of the survival model. We then propose extrapolation methods based on flexible parametric polyhazard models which can naturally accommodate diverse shapes, including non-proportional hazards and crossing survival curves, while typically maintaining a natural interpretation. We estimate the mean survival and related estimands in three cases, namely breast cancer, cardiac arrhythmia and advanced melanoma. Specifically, we evaluate the survival disadvantage of triple-negative breast cancer cases, the efficacy of combining immunotherapy with mRNA cancer therapeutics for melanoma treatment and the suitability of implantable cardioverter defibrilators for cardiac arrhythmia. The latter is conducted in a competing risks context illustrating how working on the cause-specific hazard alone minimizes potential instability. The results suggest that the proposed approach offers a flexible, interpretable and robust approach when survival extrapolation is required.
@@ 1 Introduction
Extrapolation beyond the observed survival curve can be dangerous and should not be made casually. However, practical applications necessitate such extrapolation, especially in health economic modeling and related applications of survival analysis. This is sometimes performed by fitting naive parametric models and simply projecting them to the future. Alternatively, solutions based upon the restricted mean survival (RMS) have been proposed, but from the health economic perspective these give an accurate answer to the wrong problem.
@@
Anchoring against external data has been recognised as a way to build towards robust ways of extrapolation to the future (see Rutherford et al. 2020), either from a Bayesian perspective (e.g. Demiris and Sharples 2006) or using a frequentist approach. This is essentially being proposed because it offers an implicit bias-variance trade-off in the unseen data since utilising long term information minimises the chance of wild extrapolation, albeit with the possibility of adding bias. Since this trade-off is concerned with future data, this represents a non-standard statistical problem. See Jackson et al. (2017), Bullement et al. (2023) and Latimer and Adler (2022) for recent reviews on survival extrapolation by anchoring against external data. In some recent developments, Gallacher et al. (2021) discuss problems regarding the goodness-of-fit statistics to select a model for extrapolation, while Che et al. (2023) present a method that gradually blends the predicted survival of the disease population with the survival of the general population. Guyot et al. (2017) use splines to extrapolate survival of cancer patients, while the practicality of adopting this line of research has improved substantially by the approach of Jackson (2023) who provides an R package for flexible and practical extrapolation using M-splines.
@@
Our methodology is motivated from and tested on three case studies regarding breast cancer, metastatic skin melanoma and cardiac arrhythmia. The breast cancer dataset contains long follow up and therefore facilitates the thorough assessment of the predictive ability of the proposed approach based on suitable information fractions. We use individual information on particular genes to classify individuals as triple negative or not and evaluate their life expectancy in a challenging scenario of crossing survival curves. The current developments regarding mRNA therapeutics is tackled in melanoma where we estimate the advantage when the mRNA vaccine is received along with pembrolizumab versus those receiving pembrolizumab immunotherapy alone. Our third use case is concerned with an arrhythmia dataset where we estimate the LYG for patients that receive an implantable cardioverter defibrilator (ICD) compared to a group receiving anti arrhythmia drugs (AAD). In doing so, we extend the poly-hazard models with the introduction of suitable change-points. The implementation of proposed methods has been carried out in the R language and the relevant Stan code is made freely available on Github.
@@ 2.2 Advanced Melanoma
Melanoma refers to a type of skin cancer at metastatic stage. Their treatment has been revolutionalized due to immunotherapy such as pembrolizumab; see Villani et al. (2022) for a discussion of the currently available treatment options. Perhaps the most exciting new cancer treatment is based on mRNA technology and a number of clinical trials are currently being deployed, mostly on patients with resected melanoma at high risk of recurrence and is being used as an adjuvant therapy in combination with adjuvant pembrolizumab. Khattak et al. (2023) have published the first such trial where the hazard ratio for recurrence-free survival was 0.561 for patients receiving ‘mRNA-4157 (V940) + pembro’ against pembro alone, while no estimates for overall survival were reported. Preliminary survival estimates of these groups are thus desirable. We digitize the published survival curve (see the on-line Appendix A) from Hamid et al. (2019) to obtain event times of patients receiving pembro. We then use the estimated hazard ratio as the basis for the ‘mRNA+pembro’ curve. In the on-line Appendix F, we gather the results from a more recent clinical trial (Robert et al. 2023).
@@ 4.1 Breast Cancer
The breast cancer dataset has long follow up and hence a nearly complete survival curve, thus facilitating repeated testing of the trained models against the ‘truth’ in the form of the observed data. We use various levels of the information fraction (Demets and Lan 1994) as a metric. For example, 80% of information corresponds to 49% survival probability. The best model for this dataset was a joint Bi-Weibull for the disease group and the external population. We impose the constraint that the first components of each Bi-Weibull model will be proportional to each other and the second components will be equal. Thus, we generate age-sex matched data from the external population $y_{1:n_{p}}$ and we assume each sample $i=1,...,n_{p}$ has hazard $h_{p}(y_{i};\alpha_{1},\alpha_{2},\lambda_{1},\lambda_{2})=h_{p}^{1}(y_{i};\alpha_{1},\lambda_{1})+h_{p}^{2}(y_{i};\alpha_{2},\lambda_{2})$. The hazard form of each component $m$ is Weibull with shape and rate parameters $\alpha_{m}$ and $\lambda_{m}$ respectively, i.e. $h_{p}^{m}(y_{i};\alpha_{m},\lambda_{m})=\lambda_{m}\cdot\alpha_{m}\cdot y_{i}^{\alpha_{m}-1}$. The survival function is $S(y_{i};\alpha_{1},\alpha_{2},\lambda_{1},\lambda_{2})=S_{p}^{1}(y_{i};\alpha_{1},\lambda_{1})\cdot S_{p}^{2}(y_{i};\alpha_{1},\lambda_{1})$, where each component takes the form $S_{p}^{m}(y_{i};\alpha_{m},\lambda_{m})=\exp(-\lambda_{m}\cdot y_{i}^{\alpha_{m}})$. The likelihood is given by equation (2). For the disease data $t_{1:n_{d}}$, we also assume a Bi-Weibull model, but every sample $i=1,...,n_{d}$ has hazard $h_{d}(t_{i};\alpha_{1},\alpha_{2},\lambda_{1},\lambda_{2})=C\cdot h_{p}^{1}(t_{i};\alpha_{1},\lambda_{1})+h_{p}^{2}(t_{i};\alpha_{2},\lambda_{2})$ and survival $S(t_{i};\alpha_{1},\alpha_{2},\lambda_{1},\lambda_{2})=S_{p}^{1}(t_{i};\alpha_{1},\lambda_{1})^{C}\cdot S_{p}^{2}(t_{i};\alpha_{1},\lambda_{1})$, where $C>0$. The likelihood is given by (1) and the complete likelihood by (3).
@@ 4.1.1 Subgroup Analysis for precision medicine
4.2 The effect of mRNA therapeutic on Melanoma
@@
This treatment is used as an adjuvant therapy in combination with adjuvant pembrolizumab in patients with resected melanoma at high risk of recurrence. The best fitted model for the overall survival in the advanced melanoma dataset was a joint Tri-Weibull where the first components of the disease and external groups are proportional to each other while the third components are equal. Thus, we use $m=3$ components: $h_{p}(y_{i};\alpha_{1},\alpha_{2},\alpha_{3},\lambda_{1},\lambda_{2},\lambda_{3})=h_{p}^{1}(y_{i};\alpha_{1},\lambda_{1})+h_{p}^{2}(y_{i};\alpha_{2},\lambda_{2})+h_{p}^{3}(y_{i};\alpha_{3},\lambda_{3})$ and the hazard of the pembrolizumab group is given by $h_{pembro}(t_{i};\alpha_{1},\alpha_{d},\alpha_{3},\lambda_{1},\lambda_{d},\lambda_{3})=C\cdot h_{p}^{1}(t_{i};\alpha_{1},\lambda_{1})+h_{d}^{2}(t_{i};\alpha_{d},\lambda_{d})+h_{p}(t_{i};\alpha_{3},\lambda_{3})$. In this example we select the ‘constant difference’ method and use the last 5 points (months) to write $h_{pembro}(t^{\prime})=D_{pembro}+h_{p}(t^{\prime})$ for a future time $t^{\prime}$. The mRNA group of patients is derived via the proportional hazards assumption $h_{mRNA}(t)=0.561\cdot h_{pembro}(t)$ (see Khattak et al. 2023) where we assume that the reported hazard ratio (HR) may reasonably approximate the HR for overall survival. Extrapolation follows using $h_{mRNA}(t^{\prime})=D_{mRNA}+h_{p}(t^{\prime})$ for future times $t^{\prime}$.
Figure 3 depicts the extrapolated survival curves and hazard functions. Calculating the area between the mRNA and pembrolizumab curves estimates that the adjuvant mRNA therapeutic offers on average an additional 43.69 months of life (see right plot of Figure 4). One may broadly interpret this as patients with advanced melanoma may gain 3.64 years of life if the mRNA therapeutic is added to their treatment. In the absence of HR estimates for overall survival we used the corresponding HR estimate for recurrence free survival and suitable sensitivity analyses ought to be adopted if these estimates are to be used for decision support/making.
@@ 5 Discussion
Fitting independent poly-hazard models to the external population and the disease groups can give seemingly good fit during the follow-up period at the cost of unstable and/or unreliable extrapolation. This article uses methods that can avoid such erratic behaviour by transferring appropriate information from data external to the disease group. The external population is based upon mortality projections as opposed to past data and the evidence is synthesized in a Bayesian context. We found that modeling on the hazard scale can identify incoherent fit more efficiently compared to the survival and cumulative hazard scale. We employ an extension of the standard poly-hazard models introducing change-points, offering added flexibility.
@@
A review of tests for crossing survival curves may be found at Li et al. (2015). Our work is concerned with a more involved issue, not simply testing for equality, but fully modeling the survival curves and the calculation of their difference, as necessitated in health economic evaluations. To do so, we transfer long term evidence to the current context thus facilitating relatively stable extrapolation. Alternatives to poly-hazard models include splines (Jackson 2023) and fractional polynomials (Royston and Parmar 2002). Both approaches offer nearly arbitrary flexibility at the cost of typically limited interpretability of their parameters. Perhaps more importantly, poly-hazard models offer a natural balance between efficient use of the data and undesirable carry-over effects.
@@
This work was motivated by 3 use cases that include a crossing survival problem, an emerging cancer therapeutic and a competing risks example. The unified approach presented here naturally caters for survival extrapolation in these cases. We estimate that women with breast cancer lose on average 10.17 years compared to the general population and we also estimate the additional deficit of the triple negative sub-population. The mRNA-4157 (V940) therapeutic in combination with pembrolizumab is currently being assessed for treatment of melanoma and by incorporating the early trial evidence within our framework we estimate that an average of 3.64 extra years of life may be gained compared to pembrolizumab alone. The cardiac arrhythmia estimates mortality reduction when switching from AAD to ICD treatment by retaining a common hazard on death from other causes.
@@
The results are assessed under five extrapolation methods. The baseline method performs well in all our examples by appropriately modelling the bond between the disease and external populations. The constant difference and ratio methods build upon the findings of the follow-up period and extrapolate the last observed relationship between the two populations into the future. We further amend these two methods by applying them to the components of interest alone and obtain the cause-specific analogues. The proposed methodology does require individual-level data. This problem can be resolved by digitizing published Kaplan-Meier curves as illustrated in the melanoma and cardiac arrhythmia data. Still, some inaccuracies may result and care is required in this aspect. In addition, expert opinion may be elicited as done by Benaglia et al. (2015) for the age-specific proportions of arrhythmia.
@@
Survival analysis is also used in non-medical application areas of statistics, including industrial, actuarial and reliability functions where the idea of transferring information from an external source can be fruitful if extrapolation is required. Further development of the piecewise poly-hazard models may be considered by training them based upon alternative combinations of the disease and external groups. The predictive ability of the models developed in this article has been informally assessed, since in reality it is dictated by the performance on future/unseen data. One alternative is to characterize the model’s properties using an appropriate loss function on the combined clinical and registry data and this represents the subject of current work.
Acknowledgements
The authors are grateful to Dr. Pippa Corrie for comments on the melanoma example.
@@ A. The details of KM digitization
B. Some comments on typical extrapolations
@@ C. Mortality projections
D. Precision breast cancer results
Results on the extrapolation of breast cancer data related with the 3N and n3N groups.